OpenStudy (anonymous):

please help solve 5 x^2+ 3 x + xy = 4 and y( 4 ) = -22, find y'( 4 ) by implicit differentiation.

OpenStudy (amistre64):

5 x^2+ 3 x + xy = 4 10x+3 +x'y + xy' = 0 10x+3 +y + xy' = 0 solve for y' and plug in a 4

OpenStudy (anonymous):

I know that you must make the equation y equal: y=-(5x^2+3x-4)/x

OpenStudy (amistre64):

implicit means dont make it equal to y ;)

OpenStudy (anonymous):

I know but to sove it it would be easier to do that

OpenStudy (anonymous):

The derivative I think would be: (10x+3)/x^2

OpenStudy (amistre64):

in this case, maybe; implicits are fairly easy. The only caveat is that you get a y in the implicit that would have to be subbed in from the explicit anyhows

OpenStudy (anonymous):

Oh forget it I don't know what implicit means

OpenStudy (amistre64):

lol, implicit means solve it as tho you solved for y; but without solving for y

OpenStudy (anonymous):

I gotta learn how to do that

OpenStudy (anonymous):

i was solving it explicit and but I just couldn't get how to work the y( 4 ) = -22 and the find y'( 4)

OpenStudy (anonymous):

implicitly**

OpenStudy (amistre64):

i got it, since y(4) = -22; just sub that in for y 10x+3 +y + xy' = 0 10(4)+3 +-22 + (4)y' = 0 40+3 -22+ (4)y' = 0 21+ (4)y' = 0 y' = -21/4

OpenStudy (anonymous):

Oh that is cool

OpenStudy (anonymous):

ahhh thank you so much, i think i was making it so much more complicated for myself and thank you for your help too rld613

OpenStudy (amistre64):

implicit is just a name, and they really should teach you that its nothing new from what your used to

OpenStudy (amistre64):

all you do is NOT throw out your derived bits; those x' and y'

OpenStudy (amistre64):

the rules for the rest of it are exactly the same

OpenStudy (amistre64):

product rule: [fg]' = f'g+fg' right? it doenst matter what the letters are: [xy]' = x'y+xy'

OpenStudy (amistre64):

since your used to deriving with respect to "x"; that x' (or rather dx/dx) = 1 and you toss it by habit

OpenStudy (anonymous):

That is true

OpenStudy (amistre64):

but ... you are perfectly content with y going to y' from the start

OpenStudy (anonymous):

I have never done it like that before

OpenStudy (amistre64):

tell me, what is the derivative of x^3?

OpenStudy (anonymous):

3x^2

OpenStudy (amistre64):

almost; i never stated with respect to x; so its actually 3x^2 x' until we KNOW what to do with the x'

OpenStudy (amistre64):

whats the derivative of 2y^2?

OpenStudy (anonymous):

4yy'

OpenStudy (amistre64):

exactly

OpenStudy (anonymous):

I don't understand why?

OpenStudy (amistre64):

the rules and processes are not changed, you just learn not to toss out the derived bits

OpenStudy (anonymous):

I don't understand why?

OpenStudy (anonymous):

I don't understand why?

OpenStudy (anonymous):

I don't understand why?

OpenStudy (anonymous):

I don't understand why?

OpenStudy (anonymous):

I don't understand why?

OpenStudy (anonymous):

I don't understand why?

OpenStudy (anonymous):

I don't understand why?

OpenStudy (amistre64):

lets try it this manner: \[\frac{d}{dx}x^3\] \[\frac{dx}{dx}3x^2\] \[1*3x^2\] \[3x^2\] right?

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

that is what you are used to doing ... and you already do it for the y to begin with: \[\frac{d}{dx}y\] \[\frac{dy}{dx}1\] \[\frac{dy}{dx}=y'\]

OpenStudy (amistre64):

or maybe: \[\frac{d}{dx}2y^2\] \[\frac{dy}{dx}4y\]

OpenStudy (anonymous):

y is it negative?

OpenStudy (amistre64):

the posted problem is negative becasue that is the slope at x=4

OpenStudy (amistre64):

it works out the same whether it was explicit or implicit

OpenStudy (anonymous):

I gotta lookt his over. hopefully i will figure it out. It as to sink in

OpenStudy (amistre64):

good luck with it :)

OpenStudy (anonymous):

thanks for ur time